# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

### v1

@@ IV-A Global Search
Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization 37, 24. These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima. Each candidate solution is evaluated using a fitness function, $F(X)$, which quantifies the quality of SRO in the atomic structure by comparing local atomic arrangements to the desired SRO characteristics.
@@ IV-B Local Search
The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations. To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Algorithm 3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel. This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained. This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
Compared to the serial version (Algorithm 2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously. This parallelization significantly enhances both scalability and performance. By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
IV-C Short-Range Order
Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution 38. The SRO is typically quantified by the Warren-Cowley parameter 9, denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$. For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
@@ IV-C Short-Range Order
IV-D Molecular Dynamics
Molecular dynamics (MD) simulations were performed using LAMMPS 39, 40, 41, employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system. The simulations encompassed system sizes ranging from 256 to 260,000 atoms, with a timestep of 2 fs in an NPT ensemble. The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies. The results of these MD simulations are discussed in detail in Section II, demonstrating the effectiveness of our method.
IV-E Incremental Calculation of SRO (Fitness)
@@
In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Eqs. 2). This recalculation becomes increasingly computationally expensive as system size grows. Fig. 5 illustrates that for larger systems, the SRO calculation can consume more than 99% of the total computational cost, highlighting the necessity for a more efficient method.
@@ IV-F1 Parallel Processing on the GPU
  • Lattice Structure Storage: The lattice structure 42 is stored in shared memory to minimize memory access time during each MC iteration. This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
IV-F2 Iterative Fitness Calculation
@@
IV-F3 Fitness Sorting and Selection
@@
IV-F4 Convergence Check and Termination

### v2

@@ 4.1 Global Search
The optimization of atomic configurations in High-Entropy Alloys (HEAs) presents a complex challenge due to the vast solution space. As illustrated by Fig. 4, our global search framework addresses this challenge through an iterative approach that combines systematic exploration with targeted refinement.
@@
Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization [36, 23]. These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima. Each candidate solution is evaluated using a fitness function, $F(X)$, which quantifies the quality of SRO in the atomic structure by comparing local atomic arrangements to the desired SRO characteristics.
@@ 4.2 Local Search
The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations. To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Alg. 3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel. This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained. This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
Compared to the serial version (Alg. 2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously. This parallelization significantly enhances both scalability and performance. By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
4.3 Short-Range Order
Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution [37]. The SRO is typically quantified by the Warren-Cowley parameter [9], denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$. For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
@@ 4.3 Short-Range Order
4.4 Molecular Dynamics
Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40], employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system. The simulations encompassed system sizes ranging from 256 to 256,000 atoms, with a timestep of 2 fs in an NPT ensemble. The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies. The results of these MD simulations are discussed in detail in section 2, demonstrating the effectiveness of our method.
4.5 Incremental Calculation of SRO (Fitness)
@@
In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Equation 3). This recalculation becomes increasingly computationally expensive as system size grows. Fig. 5 illustrates that for larger systems, the SRO calculation can consume more than 99% of the total computational cost, highlighting the necessity for a more efficient method.
@@ 4.6.1 Parallel Processing on the GPU
  • 2.
    Lattice Structure Storage: The lattice structure [41] is stored in shared memory to minimize memory access time during each MC iteration. This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
4.6.2 Iterative Fitness Calculation
@@
4.6.3 Fitness Sorting and Selection
@@
4.6.4 Convergence Check and Termination
@@ 6 Code Availability
Appendix A Supplementary Data
Table 2: Performance comparison between PyHEA and existing methods under parallel computing conditions. Both PyHEA and SCRAPs utilized 24 CPU cores (Intel Xeon Gold 6132 @ 2.60GHz) for parallel execution, with additional GPU (single NVIDIA 4090) results shown for PyHEA. The table presents computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures of varying sizes (#At) and species numbers (#Sp). All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks (depth 10), and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc). Short-range order was optimized over 3 coordination shells. Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
  Type #Spa #Atb ATAT SCRAPs PyHEA PyHEA Speedupc
Parallel Parallel GPU CPU GPU
  bcc 3 54 >1.4kd 0.14 0.14 .03 1.0[10k] 4.67[47k]
bcc 4 128 >10kd 0.34 0.19 .03 1.79[53k] 11.33[333k]
bcc 4 432 1.26 0.28 .03 4.5 42
bcc 4 1024 3.18 0.32 .03 9.94 106
bcc 4 2048 7.02 0.36 .04 19.5 175.5
bcc 5 2048 7.27 0.48 .04 15.2 181.8
bcc 5 6.8k 47.4 0.71 .05 66.8 948
bcc 6 28k 2.05 .09
fcc 3 108 0.63 0.17 .01 3.71 63
fcc 4 256 2.24 0.27 .01 8.3 224
fcc 5 500 5.81 0.43 .01 13.5 581
fcc 5 4k 64.3 0.65 .01 99.0 6.4k
fcc 4 256k 61.1 .58
  • a
    Notation: #Sp = number of elemental species.
  • b
    Notation: #At = total number of atoms in the system.
  • c
    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
  • d
    ATAT data from Ref. [23]. All times in minutes.
Figure 6: Warren–Cowley SRO parameter matrices for a multi-principal-element alloy (Fe, Mn, Cr, Co) at three different temperatures: (a) 1373 K, (b) 873 K, and (c) 873 K. The color scale ranges from blue (positive values) to red (negative values), indicating a preference for certain atom types to cluster together (positive SRO) or to avoid each other (negative SRO). At high temperatures (1373 K), the SRO values are near zero, reflecting a more random distribution, whereas at lower temperatures (873 K), the emergence of pronounced SRO patterns reveals stronger local atomic ordering.